A concentric spherical capacitor is charged to -Q on the outer sphere and +Q on the inner sphere. The inner sphere is solid, with radius=a, while the outer sphere is hollow, with radius=b. Between the two spheres is a vaccum.

a) Find the potential difference between r=a and r=b and thus find the capacitance of this system directly.

b) Assuming the outer shell is grounded (ie, at zero potential) use A concentric spherical capacitor is charged to -Q to find the capacirtance.

c) Using the energy density in the electric field find the capacitance for this pair of conductors.
Answer #1

PART A)

Potential difference is given by

V = KQ(\frac{1}{a}-\frac{1}{b})

now the capacitance is given by

C = \frac{Q}{V}

C = \frac{Q}{KQ(\frac{1}{a}-\frac{1}{b})}

C = \frac{4\pi \epsilon_0}{(\frac{1}{a}-\frac{1}{b})}

PART B)

Energy is given by

U = \frac{1}{2}CV^2

C = \frac{2U}{V^2}

PART C)

U = \frac{1}{2}CV^2

U = \frac{1}{2}*\frac{\epsilon_0 A}{d}*{E^2.d^2}

U = \frac{1}{2}*{\epsilon_0 E^2 A.d}

U = \frac{1}{2}*{\epsilon_0 E^2 }*Volume

so energy density is

U_v = \frac{1}{2}*{\epsilon_0 E^2 }